# Giant and anisotropic magnetostriction in $\beta$-O2 at 110 T

### v1

@@ Abstract
Figure 4: (a) Intermolecular potential of the dimer (O2)2 in the parallel geometry calculated in Ref. [28] as a function of total spin $S_{(\rm{O}_{2})_{2}}=0$, 1, and 2. We plotted the states $S_{(\rm{O}_{2})_{2}}=1$ and $S_{(\rm{O}_{2})_{2}}=2$ with the Zeeman effect at 110 T. (b) Magnification of a part of (a). (c) Exchange constant $J$ of spins between neighboring molecules as a function of $R$ calculated [29, 30] and deduced from a high pressure study [31]. (d) Magnification of a part of (c).
A spin-dependent intermolecular potential calculated for the dimer (O2)2 in parallel geometry shows also qualitative agreement with the above discussion [28]. In Figs. 4(a) and 4(b) shows the intermolecular potential with a variation of $S_{(\rm{O}_{2})_{2}}=0$, 1, and 2, where the data points are taken from Ref. [28] and the curves are our fit to the Morse potential $M(R)=d\left(1-\left(1-\exp\left(-\left(R-R_{0}\right)/a\right)\right)^{2}\right)$ with $d$, $a$, and $R_{0}$ being the depth, the inverted width, and the equilibrium position of the potential, respectively. As shown in Fig. 4(b), $S_{(\rm{O}_{2})_{2}}=0$ has the lowest energy with the first excited state and the second excited state being $S_{(\rm{O}_{2})_{2}}=1$ and $S_{(\rm{O}_{2})_{2}}=2$, agreeing with the antiferromagnetic spin coupling between O2 molecules in parallel geometry. Note also that the value of equilibriated $R$ increases with increasing $S_{(\rm{O}_{2})_{2}}$ as indicated by the vertical dotted lines. As the dashed lines indicate, the spin triplet and spin quintet state become more stable than the spin singlet state due to the Zeeman effect at 110 T. This trend in the intermolecular potential qualitatively explains the observed trend in the O2 crystal. The dependence of the exchange constant $J$ on the intermolecular distance of O2 is estimated by calculations [29, 30] and a high-pressure experiment of solid O2 [31] as shown in Figs. 4(c) and 4(d). These all show that $J$ is decreasing with the increasing $R$ in good agreement with our model. $\Delta J/J$ varies from -3 to -5 K/bond depending on the curves in Fig. 4(d), which amounts to -9 to -15 K/O2, which is a sufficient value for the lattice to move for 1 % considering the energy scale in the intermolecular potentials in Figs. 4(a) and 4(b).
@@
The qualitative argument above with the dimer (O2)2 is further strengthened by our first-principles calculation of the exchange constants in the solid $\beta$ phase. Here, we denote the nearest-neighbor exchange interaction in the $ab$ plane direction and out-of-plane directions as $J_{1}$ and $J_{2}$, respectively. The calculation uses the vdW density functional theory and the tight-binding model calculation. The calculated result of $J_{1}$(0 T) and $J_{1}$(110 T) are 47.35 and 41.8 K/bond, where $\Delta J_{1}=-5.55$ K is an 11.8 % reduction of the AFM interaction, being in good agreement with the observed expansion of the lattice constant in the $ab$ plane with increasing ferromagnetic correlation. The calculated result of $J_{2}$(0 T) and $J_{2}$(110 T) are 6.5 and 7.4 K/bond, where $\Delta J_{2}=+0.9$ K is an 14.3 % increase of the AFM interaction. $\Delta J_{2}$ is less significant than $\Delta J_{1}$ and the sign counters to the observed lattice change, indicating that the shrinkage in $c$ axis at 110 T is driven by the compensation of the lattice volume expansion due to expanding $ab$ plane rather than the exchange striction mechanism with $J_{2}$.
The significant magnetostriction may be due to the large fluctuation of the spins on the triangular lattice, which is strongly coupled to the crystal lattice. The temperature-induced lattice change in the $ab$ plane of $\beta$-O2 is as significant as 2 % with $\Delta T=20$ K, indicating that phonon and magnetic contribution to the lattice is crucial in $\beta$-O2. The uniqueness of the lattice of $\beta$-O2 is highlighted by the fact that the lattice change is much more minor in $\alpha$-O2 where the spins become rigid with the Néel order. The significance of the lattice change is consistent with previous XRD studies on the $\alpha$ phase of solid O2 at 5 T [32] and 28 T [33]. First, a large $\Delta V/V=1$ is reported at 5 T at 7 K [32], later confirmed as an extrinsic effect by Ref. [33]. Ref. [33] further put a constrain that $\Delta d/d<1\times 10^{-4}$ at 5 T and $\Delta d/d<2\times 10^{-3}$ at 25 T. If we assume the quadratic dependence $\Delta d/d\simeq pB^{2}$ with typical values of $\Delta d/d=1$ and $B=100$ T, we obtain $\Delta d/d\simeq 2.5\times 10^{-5}$ at 5 T, $6.3\times 10^{-4}$ at 25 T by an interpolation, which are consistent with the previous XRD study up to 28 T [33] and the present result.
The present lattice change occurs in the unit cell without phase transition, which is considered the greatest ever reported to our best knowledge. Larger magnetostrictions are observed in Heusller alloy [34] and manganite systems [35], where the domain’s reorientation and the phase transition are the sources of the significant lattice change. Lastly, we should observe the complete lattice transformation to $\theta$-O2 [8] above 120 T. The clarification of the crystal structure of $\theta$-O2 is, without a doubt, the most crucial goal of our study to follow. We are currently evaluating methods to increase the maximum magnetic field that PINK-02 generates.
@@ Acknowledgements.
A fruitful discussion with J. Nasu is acknowledged. The experiment was carried out with the approval of JASRI (Proposal No. 2024A8010 and 2024B8046). The authors would like to acknowledge the support from the technical staff of the SACLA facility. They also appreciate the support from the SACLA/SPring-8 Basic Development Program (2021 - 2024). This work is partially supported by JST FOREST program No. JPMJFR [^2] W and No. JPMJFR2037, JSPS Grant-in-Aid for Scientific Research on Innovative Areas (A) (1000 T Science) 23H04861, 23H04859, 24H01633, and Grant-in-Aid for Scientific Research (B) 23H01121, and Grant-in-Aid for Scientific Research 24K21043, and MEXT LEADER program No. JPMXS0320210021.
Supplementary video
Please see the supplementary video of PINK-02 generating 110 T at BL3 in SACLA 13 times successfully in a row (37 seconds). The same video is also uploaded to YouTube.

### v2

@@ Abstract
Figure 4: (a) Intermolecular potential of the dimer (O2)2 calculated as a function of the inter-molecular distance $R$ in H geometry with a variation of the total spin $S_{(\rm{O}_{2})_{2}}=0$, 1, and 2 [32]. (b) A magnification of (a). (c) Exchange constant $J_{1}$ of between molecular spins as a function of $R$ deduced from the high-pressure experiment [33]. (d) A magnification of (c).
[…]
@@
The qualitative argument above with the dimer (O2)2 is further strengthened by our first-principles calculation of the exchange constants in the solid $\beta$ phase with the structure of 0 T and the deformed unit cell at 110 T obtained in the present study. Here, we denote the nearest-neighbor exchange interaction in the $ab$ plane direction and in the interlayer ($c$) direction as $J_{1}$ and $J_{2}$, respectively. The calculation employs the vdW density functional theory and the tight-binding model calculation. The calculated result of $J_{1}$(0 T) and $J_{1}$(110 T) are 47.35 and 41.8 K/bond, where $\Delta J_{1}=-5.55$ K is an 11.8 % reduction of the AFM interaction, being in good agreement with the reported results on the dimer (O2)2 as shown in Fig. 4(d) and the exchange striction mechanism driving the expantion in the $ab$ plane. The calculated result of $J_{2}$(0 T) and $J_{2}$(110 T) are 6.5 and 7.4 K/bond, where $\Delta J_{2}=+0.9$ K is an 14.3 % increase of the AFM interaction. $\Delta J_{2}$ is less significant than $\Delta J_{1}$ and the sign counters to the observed lattice change. This indicates that the shrinkage in $c$ axis at 110 T is not driven by the exchange striction mechanism with $\Delta J_{2}$, which accords with the previous notion that the inter-layer coupling of spins is negligible due to the small hybridization of the $\pi$ orbitals [35]. Another possible mechanism for the shrinkage in $c$ axis is the conservation effect of the lattice volume expansion with the expanding $ab$ plane, which is not conclusive.
In conclusion, we obtained powder x-ray diffraction data of $\beta$-O2 at 110 T by establishing the novel portable 100 T technique PINK-02 compatible with XFEL. An anisotropic lattice deformation of up to 1 % was observed, which arises from the two-dimensionality of the spin system and the competition between exchange and van der Waals interactions. Our method opens the way to exploring spin–lattice coupling phenomena above 100 T. The clarification of the crystal structure of $\theta$-O2 is, without a doubt, the most crucial goal of our study to follow.
@@ Acknowledgements.
Fruitful discussion with J. Nasu is acknowledged. The experiment was conducted with the approval of JASRI (Proposal Nos. 2024A8010 and 2024B8046). The authors would like to acknowledge the support from the technical staff of the SACLA facility. PINK-02 is established with the support of the SACLA/SPring-8 Basic Development Program (2021 - 2024). This work is supported by the JST FOREST (Program Nos. JPMJFR [^2] W and JPMJFR2037), JSPS Grant-in-Aid for Scientific Research on Innovative Areas (A) (1000 T Science) 23H04861, 23H04859, 24H01633, Grant-in-Aid for Scientific Research (B) 23H01121, Grant-in-Aid for Scientific Research 24K21043, and MEXT LEADER program No. JPMXS0320210021.
Data Availability
The data that support the findings of this article are openly available  [^4] https://doi.org/10.5281/zenodo.17166637.